Sachs Brands defined benefit pension plan specifies annual retirement benefits equal to: 1.4% x service years x final years salary, payable at the end of each year. Angela Davenport was hired by Sachs at the beginning of 2002 and is expected to retire at the end of 2036 after 35 years service. Her retirement is expected to span 18 years. Davenports salary is $92,000 at the end of 2016 and the companys actuary projects her salary to be $290,000 at retirement. The actuarys discount rate is 6%. (FV of $1, PV of $1, FVA of $1, PVA of $1, FVAD of $1 and PVAD of $1) (Use appropriate factor(s) from the tables provided.) Required 2. Estimate by the projected benefits approach the amount of Davenports annual retirement payments earned as of the end of 2016. Annual retireement payments 3. What is the companys projected benefit obligation at the end of 2016 with respect to Davenport? Projected benefit obligation 4. If no estimates are changed in the meantime, what will be the companys projected benefit obligation at the end of 2019 (three years later) with respect to Davenport? Projected benefit obligationPlease HELP ASAP!!
Answer #1

(2) Davenport's annual retirement payments at the end of 2016

Service years = 14 years (2002-2016)

final year's salary = $ 92,000

Annual retirement payments =1.4% * 14 years * 92,000= $ 18,032

(3) Projected benefit obligation at the end of 2016

Present value of the annual retirement benefits on the retirement date where n =18 years and i = 6%

= 18032 * PVIF for annuity

=18032 * 10.8276

= $195,243.30

Present value of the retirement benefits at the end of 2016

= PV of annuity on retirement date * PVIF of single cash flow when n= 20 years and i= 6%

= 195243.30 * 0.3118 = $ 60,876.86

(4) Projected benefit obligation at the end of 2019 if estimates don't change(assuming salary at the end of 2019 is $92,000)

Service years = 17 years (2002-2019)

final year's salary = $ 92,000

  Annual retirement payments =1.4% * 14 years * 92,000= $ 21,896

  

Present value of the annual retirement benefits on the retirement date where n =18 years and i = 6%

= 21896 * PVIF for annuity

=18032 * 10.8276

= $237,081.10

Present value of the retirement benefits at the end of 2019

= PV of annuity on retirement date * PVIF of single cash flow when n= 17 years and i= 6%

= 237081.10 * 0.3714 = $ 88,051.92